OpenStudy (anonymous):

How would you find the domain of 4/3x-5

OpenStudy (tkhunny):

You don't FIND the Domain, you DETERMINE the Domain. It is a thought question, not a calculation question. 1) Do you mean what you have written? You wrote: \(\dfrac{4}{3}x - 5\). The Domain is All Real Numbers. 2) Did you mean \(\dfrac{4}{3x-5}\)? That is a different Domain. What makes the Denominator zero (0)? 3) You MUST see the difference between what you have written and what you intended.

OpenStudy (anonymous):

The second one

OpenStudy (tkhunny):

Not a good answer. Please respond to my point #3. After that, please respond the question in Point #2. You get to show YOUR work.

OpenStudy (anonymous):

Yes I see how the question could be written or interpreted two different ways. I'm asking if I just put the bottom equal to zero, and what do I have to do with the top number. And if you're going to put attitude in your answer please don't help me again

OpenStudy (tkhunny):

Okay, then you are not understanding. It CANNOT be interpreted two different ways. 4/3x-5 can ONLY be interpreted \(\dfrac{4}{3}x-5\) If you had used parentheses, 4/(3x-5) is very clearly, and ONLY, \(\dfrac{4}{3x-5}\). As stated in point #2, where is the denominator zero (0)? That is all this one takes. Remember your Order of Operations. Things need to make sense.

OpenStudy (anonymous):

Okay, my bad for not adding parentheses, thanks for the help